The Nelson Company has $1,485,000 in current assets and $550,000 in current liabilities. Its initial inventory level is $440,000, and it will raise funds as additional notes payable and use them to increase inventory.

  1. How much can Nelson's short-term debt (notes payable) increase without pushing its current ratio below 1.8? Round your answer to the nearest cent.

    $  


  2. What will be the firm's quick ratio after Nelson has raised the maximum amount of short-term funds? Round your answer to two decimal places.
Answer #1

1.
Let x be the increase in notes payable
(1485000+x)/(550000+x)>=1.8
=>x<=617348.65


2.
Quick ratio=(1485000-440000)/(550000+617348.65)=0.895190995